# Boundary vertex algebras for 3d $\mathcal{N}=4$ rank-0 SCFTs

### v2

@@ 4.1.2 Deformation to the $B$-twist
In order to deform to the $B$ twist we first need to set $\nu\to 1$; this ensures the operator $\phi^{2}V_{-1}$ has spin $1$ and has the effect of making all of the above generators spin $1$. The $B$-twist is then realized by introducing the superpotential $W=-\tfrac{1}{2}\phi^{2}V_{-1}$. Note that this superpotential explicitly breaks the topological flavor symmetry.
@@
$\overline{\lambda}(z)\overline{\lambda}(w)\sim\frac{\partial_{\phi}^{2}W(w)}{z-w}\quad\rightsquigarrow\quad\theta_{-}(z)\theta_{-}(w)\sim\frac{-f(w)}{z-w}\ .$ (51)
The remaining OPEs then uniquely determined by associativity:
$\theta_{+}(z)\theta_{-}(w)\sim\frac{1}{(z-w)^{2}}+\frac{h(w)}{z-w}\ ,\qquad\theta_{+}(z)\theta_{+}(w)\sim\frac{e(w)}{z-w}\ .$ (52)
@@ 4.3 Modular transformations of supercharacters
In this subsection we turn to the modular transformation properties of the modules found above; as with fusion, these can be derived from those of the ${\cal U}=L_{1}(\mathfrak{sl}(2))\otimes M(3,5)$ subalgebra, cf. Creutzig:2017anl . Note that the supercharacters of a module and its parity shift agree with one another up to a factor of $-1$, $\chi[M]=-\chi[\Pi M]$, and hence are linearly dependent of one another. Correspondingly, we will restrict our attention to the modules $\mathbf{1}$ and $\mathbf{M}$ [^4] We could equivalently consider $\mathbf{1}$ and $\Pi\mathbf{M}$ by conjugating the following $S$- and $T$-matrices by the diagonal matrix $\textrm{diag}(1,-1)$, corresponding to the change of basis from $\{\chi[\mathbf{1}],\chi[\mathbf{M}]\}$ to $\{\chi[\mathbf{1}],\chi[\Pi\mathbf{M}]\}$.
@@
This conjugation is consistent with the fact that the $y\to 1$ limit of the supercharacters of $\mathbf{1}$ and $\Pi\mathbf{M}$ precisely agree with those of the $M(2,5)$ modules $V^{(2,5)}_{1,2}$ and $V^{(2,5)}_{1,1}$, respectively. [^5] Our $S$-matrix agrees with the one appearing in Eq. (3.11) of Gang:2021hrd , but the $T$-matrix is slightly different: $T_{GKLSY}=e^{-2\pi i/60}T^{-1}\qquad S_{GKLSY}=S$ The matrices $T_{GKLSY}$ and $S_{GKLSY}$ do not satisfy the relation $(ST)^{3}=1$ of $PSL(2,\mathbb{Z})$ and $SL(2,\mathbb{Z})$.
@@ 4.4 A boundary condition for $M(2,5)$
5 Higher rank VOAs
A first-principles derivation of a boundary VOA ${\cal V}_{r}$ for the higher-rank theory ${\cal T}_{r}$ is much harder to implement than the rank 1 case. In brief, this is due to the form of the deformations proposed by GKS . Instead, we present indirect evidence for the following proposal: the $B$ twist of ${\cal T}_{r}$ admits
@@ 5.1 A deformable boundary condition for ${\cal T}_{r}$
$I\!\!I(y_{i};q)=(q;q)_{\infty}^{-r}\sum_{\vec{\mathfrak{m}}\in\mathbb{Z}^{r}}q^{\scriptsize{\frac{1}{2}}\vec{\mathfrak{m}}^{T}K_{r}\vec{\mathfrak{m}}}\vec{y}\,{}^{K_{r}\vec{\mathfrak{m}}}(y_{1}^{-1}q^{1-\mathfrak{m}_{1}};q)_{\infty}\dots(y_{r}^{-1}q^{1-\mathfrak{m}_{r}};q)_{\infty}\ .$ (80)
@@ 5.1.1 Superpotential constraints
As described in Section 3, the allowed boundary conditions are constrained by the fact that they must be compatible with the bulk superpotential. The superpotential relevant for the $B$ twist contains two parts. The first part comes from the bare monopoles $V_{\mathfrak{m}_{1}}$, $\dots$, $V_{\mathfrak{m}_{r-1}}$ which are argued by GKS to be necessary for the IR theory to exhibit an enhancement to ${\cal N}=4$ supersymmetry. The second part is given by the dressed monopole operator $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$ and is used to deformation (the $HT$ twist of) this theory to the $B$ twist of the IR SCFT.
The first constraints on the boundary condition we consider are those coming from the operator $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$. One constraint says that this operator must vanish when brought to the boundary: Dirichlet boundary conditions will break SUSY unless the superpotential vanishes at the boundary, cf. Section 2.3 of DGP , and this breaking is related to the boundary condition being deformable CG ; BLS . This term in the superpotential suggests we should require that at least one of the scalars vanishes on the boundary. On the other hand, the superpotential should not vanish to a very high order. As described in CDG , a non-vanishing $\partial_{\phi}W|$ leads to a differential on the fermions, a non-vanishing $\partial^{2}_{\phi}W|$ leads to a singular OPE, and more generally non-vanishing components of the tensor of $n$th-order derivatives $\partial_{\phi}^{n}W|$ may lead to higher $L_{\infty}$ operations on the boundary vertex algebra. To ensure that the superpotential $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$ induces a suitable singular OPE and no higher operations, we set $r-1$ of the scalars to non-zero values, with the remaining scalar forced to vanish on the boundary. [^8] A priori, the fact that the higher-order derivatives of $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$ do not vanish when we give $r-1$ of the scalars generic boundary values could lead to higher operations. Thankfully, this possibility is mitigated by the appearance of a differential that trivializes all but $1$ boundary fermion. Schematically, the differential takes the form $QB=\mu|=\phi|\overline{\lambda}$ where $\mu$ is the current generating the $U(1)$ action on the chiral multiplet; this does not arise when $\phi|=0$, but when $\phi|\neq 0$ it implements the breaking of the $U(1)_{\partial}$ symmetry as the generating current $B$ is no longer $Q$-closed.
@@
In summary, we propose that the boundary condition Dir${}^{(r)}$ defined by
@@
With the boundary condition Dir${}^{(r)}$ in hand, we now turn to the algebra of boundary local operators. Unlike the rank 1 case, we are unable to determine the algebra of local operators directly. Instead, we present some indirect evidence supporting the proposal that this algebra of local operators realizes the simple affine VOA $L_{r}(\mathfrak{osp}(1|2))$.
Our main piece of evidence for this proposal comes from comparing half-indices and (super)characters of the simple modules of $L_{r}(\mathfrak{osp}(1|2))$, cf. Creutzig:2018ogj . The half-index counting local operators on Dir${}^{(r)}$ is given by
@@
We note that the bare monopole $V_{(-1,0,\dots,0)}|$ survives on Dir${}^{(r)}$ due to its negative magnetic charge and has the necessary quantum numbers ($B_{r}$-charge $-2$ and spin $1$) to be identified with the bosonic generator $f(z)$. Moreover, the term $\phi_{1}^{2}\dots\phi_{r}^{2}V_{(-1,0,\dots 0)}$ in the superpotential introduces a non-trivial OPE between the $r$th fermion $\overline{\lambda}=\overline{\lambda}_{r,-}|$ and itself:
@@
We can again use the analysis of GKS to predict which bulk line operators are identified with which modules for $L_{r}(\mathfrak{osp}(1|2))$. Namely, we consider Wilson lines with charges $-W^{r}_{n}$ (dressed by a suitable $R$-symmetry Wilson line). The half-index counting local operators at the junction of such a Wilson line and Dir${}^{(r)}$ is given by

### v3

@@ 4.1.2 Deformation to the $B$-twist
In order to deform to the $B$ twist we first need to set $\nu\to 1$; this ensures the operator $\phi^{2}V_{-1}$ has spin $1$ and has the effect of making all of the above generators spin $1$. The $B$-twist is then realized by introducing the superpotential $W=-\phi^{2}V_{-1}$. Note that this superpotential explicitly breaks the topological flavor symmetry.
@@
$\overline{\lambda}(z)\overline{\lambda}(w)\sim\frac{\partial_{\phi}^{2}W(w)}{z-w}\quad\rightsquigarrow\quad\theta_{-}(z)\theta_{-}(w)\sim\frac{-2f(w)}{z-w}\ .$ (49)
The remaining OPEs uniquely determined by associativity, i.e. they are uniquely determined by imposing that the commutators of the modes of these generators satisfy the Jacobi identity. For example, consider the OPE of $\theta_{+}$ and $\theta_{-}$ or, equivalently, the anti-commutator $\{\theta_{+,n},\theta_{-,m}\}$. Using the relation $\theta_{+,n}=-[e_{n-m},\theta_{-,m}]$ and $[\theta_{-,m},\theta_{-,n}]=-2f_{n+m}$, it follows that
$\{\theta_{+,n},\theta_{-,m}\}$ =$=-\{[e_{n-m},\theta_{-,m}],\theta_{-,m}\}$ (50)
=[$=[\{\theta_{-,m},\theta_{-,m}\},e_{n-m}]-\{[\theta_{-,m},e_{n-m}],\theta_{-,m}\}$
$=2[e_{n-m},f_{2m}]-\{\theta_{+,n},\theta_{-,m}\}$
from which we conclude this anti-commutator is given by
$\{\theta_{+,n},\theta_{-,m}\}=[e_{n-m},f_{2m}]=h_{n+m}+2n\delta_{n+m,0}\,,$ (51)
corresponding to the OPE
$\theta_{+}(z)\theta_{-}(w)\sim\frac{2}{(z-w)^{2}}+\frac{h(w)}{z-w}\,.$ (52)
The OPE of $\theta_{+}$ with itself can be computed similarly, leading to
$\theta_{+}(z)\theta_{+}(w)\sim\frac{2e(w)}{z-w}\ .$ (53)
@@ 4.3 Modular transformations of supercharacters
In this subsection we turn to the modular transformation properties of the modules found above; as with fusion, these can be derived from those of the ${\cal U}=L_{1}(\mathfrak{sl}(2))\otimes M(3,5)$ subalgebra, cf. Creutzig:2017anl . Note that the supercharacters of a module and its parity shift agree with one another up to a factor of $-1$, $\chi[M]=-\chi[\Pi M]$. Correspondingly, we will restrict our attention to the modules $\mathbf{1}$ and $\mathbf{M}$ [^4] We could equivalently consider $\mathbf{1}$ and $\Pi\mathbf{M}$ by conjugating the following $S$- and $T$-matrices by the diagonal matrix $\textrm{diag}(1,-1)$, corresponding to the change of basis from $\{\chi[\mathbf{1}],\chi[\mathbf{M}]\}$ to $\{\chi[\mathbf{1}],\chi[\Pi\mathbf{M}]\}$.
@@
This conjugation is consistent with the fact that the $y\to 1$ limit of the supercharacters of $\mathbf{1}$ and $\Pi\mathbf{M}$ precisely agree with those of the $M(2,5)$ modules $V^{(2,5)}_{1,2}$ and $V^{(2,5)}_{1,1}$, respectively. [^5] Our $S$-matrix agrees with the one appearing in Eq. (3.11) of Gang:2021hrd , but the $T$-matrix is slightly different: $T_{GKLSY}=e^{-2\pi i/60}T^{-1}\qquad S_{GKLSY}=S$ This difference in the $T$-matrix is arises because the $T$-matrices presented in Gang:2021hrd are only determined up to an overall phase factor; indeed, the matrices $T_{GKLSY}$ and $S_{GKLSY}$ do not satisfy $(S_{GKLSY}T_{GKLSY})^{3}=1$.
@@ 4.4 A boundary condition for $M(2,5)$
5 Higher level VOAs
A first-principles derivation of a boundary VOA ${\cal V}_{r}$ for the higher-level theory ${\cal T}_{r}$ is much harder to implement than the level 1 case. In brief, this is due to the form of the deformations proposed by GKS . Instead, we present indirect evidence for the following proposal: the $B$ twist of ${\cal T}_{r}$ admits
@@ 5.1 A deformable boundary condition for ${\cal T}_{r}$
$I\!\!I(y_{i};q)=(q;q)_{\infty}^{-r}\sum_{\vec{\mathfrak{m}}\in\mathbb{Z}^{r}}q^{\scriptstyle{\frac{1}{2}}\vec{\mathfrak{m}}^{T}K_{r}\vec{\mathfrak{m}}}\vec{y}\,{}^{K_{r}\vec{\mathfrak{m}}}(y_{1}^{-1}q^{1-\mathfrak{m}_{1}};q)_{\infty}\dots(y_{r}^{-1}q^{1-\mathfrak{m}_{r}};q)_{\infty}\ .$ (81)
@@ 5.1.1 Superpotential constraints
As described in Section 3, the allowed boundary conditions are constrained by the fact that they must be compatible with the bulk superpotential. The superpotential relevant for the $B$ twist contains two parts. The first part comes from the bare monopoles $V_{\mathfrak{m}_{1}}$, $\dots$, $V_{\mathfrak{m}_{r-1}}$ which are argued by GKS to be necessary for the IR theory to exhibit an enhancement to ${\cal N}=4$ supersymmetry. The second part is given by the dressed monopole operator $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$ and is used to deform (the $HT$ twist of) this theory to the $B$ twist of the IR SCFT.
The first constraints on the boundary condition we consider are those coming from the operator $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$. One constraint says that this operator must vanish when brought to the boundary: Dirichlet boundary conditions will break SUSY unless the superpotential vanishes at the boundary, cf. Section 2.3 of DGP , and this breaking is related to the boundary condition being deformable CG ; BLS . This term in the superpotential suggests we should require that at least one of the scalars vanishes on the boundary. On the other hand, the superpotential should not vanish to a very high order. As described in CDG , a non-vanishing $\partial_{\phi}W|$ leads to a differential on the fermions, a non-vanishing $\partial^{2}_{\phi}W|$ leads to a singular OPE, and more generally non-vanishing components of the tensor of $n$th-order derivatives $\partial_{\phi}^{n}W|$ may lead to higher $L_{\infty}$ operations on the boundary vertex algebra. To ensure that the superpotential $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$ induces a suitable singular OPE and no higher operations [^8] So long as holomorphic boundary condition is deformable, it should give rise to an equivalent (quasi-isomorphic) description of line operators in the bulk TQFT. The issue with higher operations is in the complexity required to extract the $E_{2}$ or braided-tensor structure of line operators from modules of the local operators; in a sense, we are simply seeking a strict model for the category of bulk line operators., we set $r-1$ of the scalars to non-zero values, with the remaining scalar forced to vanish on the boundary. [^9] A priori, the fact that the higher-order derivatives of $\phi_{1}^{2}\dots\phi^{2}_{r}V_{(-1,0,\dots,0)}$ do not vanish when we give $r-1$ of the scalars generic boundary values could lead to higher operations. Thankfully, this possibility is mitigated by the appearance of a differential that trivializes all but $1$ boundary fermion. Schematically, the differential takes the form $QB=\mu|=\phi|\overline{\lambda}$ where $\mu$ is the current generating the $U(1)$ action on the chiral multiplet; this does not arise when $\phi|=0$, but when $\phi|\neq 0$ it implements the breaking of the $U(1)_{\partial}$ symmetry as the generating current $B$ is no longer $Q$-closed.
@@
In summary, we propose that the boundary condition Dir(r) defined by
@@
With the boundary condition Dir(r) in hand, we now turn to the algebra of boundary local operators. Unlike the level 1 case, we are unable to determine the algebra of local operators directly. Instead, we present some indirect evidence supporting the proposal that this algebra of local operators realizes the simple affine VOA $L_{r}(\mathfrak{osp}(1|2))$.
Our main piece of evidence for this proposal comes from comparing half-indices and (super)characters of the simple modules of $L_{r}(\mathfrak{osp}(1|2))$, cf. Creutzig:2018ogj . The half-index counting local operators on Dir(r) is given by
@@
We note that the bare monopole $V_{(-1,0,\dots,0)}|$ survives on Dir(r) due to its negative magnetic charge and has the necessary quantum numbers ($B_{r}$-charge $-2$ and spin $1$) to be identified with the bosonic generator $f(z)$. Moreover, the term $\phi_{1}^{2}\dots\phi_{r}^{2}V_{(-1,0,\dots 0)}$ in the superpotential introduces a non-trivial OPE between the $r$th fermion $\overline{\lambda}=\overline{\lambda}_{r,-}|$ and itself:
@@
We can again use the analysis of GKS to predict which bulk line operators are identified with which modules for $L_{r}(\mathfrak{osp}(1|2))$. Namely, we consider Wilson lines with charges $-W^{r}_{n}$ (dressed by a suitable $R$-symmetry Wilson line). The half-index counting local operators at the junction of such a Wilson line and Dir(r) is given by